OpenStudy (anonymous):

simplify 8^2/3 x 2^-1/3

OpenStudy (anonymous):

\[8=2^{3} so... 8^{(2/3)}=2^{2}\]

OpenStudy (anonymous):

then you have 2^2 * 2^(-1/3). Add the exponents

OpenStudy (anonymous):

\[8^{\frac{2}{3}}\] means cube root of 8 squared. cube root of 8 is 2, 2 squared is 4

OpenStudy (anonymous):

2 + (-1/3) = 5/3 So the answer would be 2 ^ (5/3)

OpenStudy (anonymous):

i got lost somewhere

OpenStudy (anonymous):

in order to simplify, you have to have a common base. So by simplifying the 8, you can get everything written with a base of 2

OpenStudy (anonymous):

they are added...not multiplied?

OpenStudy (anonymous):

like satellite said, the cube root of 8 is 2, so that gives you 2^2

OpenStudy (anonymous):

well, when multiplying, you add the exponents....think x^2*x^3 = x^5

OpenStudy (anonymous):

i understand the first number, not the second 2^(-1/3)

OpenStudy (anonymous):

so, since they both have a base of 2, you can add their exponents

OpenStudy (anonymous):

ok...gotcha on the addition

OpenStudy (anonymous):

You don't have to change anything in 2^(-1/3), since it alrady has a base of 2

OpenStudy (anonymous):

how do you get 5/3 by adding -1/3 and 2

OpenStudy (anonymous):

sorry im slow lol

OpenStudy (anonymous):

get common denominators....2 = 6/3

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

-1/3 + 6/3 = 5/3

OpenStudy (anonymous):

duh moment lol

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

makes sense now

OpenStudy (anonymous):

great!

OpenStudy (anonymous):

im new to this site. how exactly does it work...students helping students?

OpenStudy (anonymous):

mostly

OpenStudy (anonymous):

i have a few more of these problems, but i want to try them myself...try at least lol

OpenStudy (anonymous):

thanks for your help